Suppose the price of Good A changes to P20.00 from P18.00. Due to price change, the quantity demand of its related good which is Good B also declines from 50 units to 40 units. Compute for the cross-price elasticity of demand and determine if what type of good and elasticity


1
Expert's answer
2021-09-09T18:13:42-0400

"Priceelasticityofdemand= =\frac{\frac{(Q_2-Q_1)}{(Q_1+Q_2)/2}}{\frac{(P_2-P_1)}{(P_2+P_1)/2}}""= \frac{40-50}{40+50}" "=\frac{ -10}{90} =\frac{-1}{9}"

"\frac{18-20}{18+20} = \frac{2}{38} = \frac{1}{19}"

"= \frac{-1}{9}÷\frac{1}{19}"

"= \frac{-1}{9} × \frac{19}{1} = \frac{19}{-9}"

Price elasticity = "-2.11"